# Pedagogical Alignment of Large Language Models

### v2

@@ 4.3 Main findings: SFT vs LHP
We used the same LHP training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant performance improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated accuracy increase of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). We observed similar increase in F1 score as well: Mistral showed increase of 0.22, 0.14, and 0.21 from DPO, IPO, and KTO, and Zephyr showed improvements of 0.42, 0.12, and 0.4 respectively. Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO.
Our results are in alignment with the findings of similar studies [27], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [28] and the Ultra-Feedback dataset [14].
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements using LHP algorithms compared to SFT.
4.4 Examining LHP through lens of perplexity
@@
The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5). By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1). With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4). This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart (Action 2), and similarly for Actions 4 and 5.
Building upon the test split (450 conversations) of the originally generated dataset, we generate an alternative opposite action (Action 2 for Action 1, and Action 5 for Action 4) for the first occurrence of Action 1 or 4 in each conversation. We then use this newly constructed Action dataset to measure the perplexity of responses from the original instruct model (Base), its SFT fine-tuned counterpart (SFT), and its LHP version (DPO) (Table 3).
Our analysis reveals that for both Mistral and Zephyr Base models, responses with Action 1 have higher perplexity than those with Action 2, and likewise for Actions 4 and 5. This suggests that the base models are more inclined to provide direct solutions than offer hints and guidance. The SFT process adjusts this behavior by increasing the perplexity of responses with Actions 2 and 5, thus making the models more likely to produce responses with Actions 1 and 4. However, SFT alone doesn’t fully optimize this behavior. As seen in the perplexity results for Mistral and Zephyr SFT models (Table 3), there is only a minimal decrease or even a slight increase in the perplexity of Action 1 and 4. This indicates that SFT, while useful, doesn’t significantly shift the model’s behavior towards providing more guidance over direct solutions.
In contrast, LHP dramatically optimizes this behavior. For the Mistral model, the perplexity of Action 1 and 4 decreases from 2.05 and 2.14 to 1.89 and 1.96 (SFT vs DPO) respectively. Similarly, for the Zephyr model, the perplexity of Action 1 and 4 decreases from 2.09 and 2.18 (SFT) to 1.97 and 2.05 (DPO) respectively. This decrease in perplexity indicates a higher likelihood of the model providing guidance over direct solutions.
Simultaneously, the perplexity of Action 2 and 5 increases and from 13.9 and 13.06 (Zephyr SFT) to 14.9 and 14.38 (Zephyr DPO). This increase in perplexity indicates a lower likelihood of the model providing direct solutions, further emphasizing the model’s shift towards providing guidance.
@@
Mistral (Base) 2.11 2.06 2.00 1.97
Mistral (SFT) 2.05 5.55 2.14 5.40
Mistral (DPO) 1.89 5.48 1.96 5.34
Zephyr (Base) 2.05 2.09 2.03 2.07
@@
Zephyr (DPO) 1.97 14.9 2.05 14.38
Table 3: Perplexity of responses. Perplexity analysis reveals that the base models of Mistral and Zephyr are more inclined to provide direct solutions (Action 2 and 5) than guidance and hints (Action 1 and 4). The SFT process partially shifts this behavior, increasing the perplexity on responses with Action 2 and 5, thus making the models more likely to produce responses with Action 1 and 4. However, the SFT process alone doesn’t fully optimize this behavior, as seen by the minimal decrease or slight increase in the perplexity of Action 1 and 4. In contrast, DPO dramatically optimizes this behavior by reducing perplexity on Action 1 and 4 while simultaneously increasing perplexity on Action 2 and 5. This shift in perplexity indicates a significant behavioral shift in the LLM, making it more aligned with the role of a tutorbot by favoring guidance over direct solutions. For more details about actions please refer to Table 1.
4.5 Effect of Beta on DPO and IPO
Figure 3: DPO and IPO performance as a function of beta. We found that classification accuracy and F1 score using both DPO and IPO were significantly influenced by beta hyper-parameter adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy and 0.56 F1 score, while a higher beta of 0.9 notably reduced accuracy and F1 score. In contrast, IPO performed best with a beta of 0.1 and increasing beta decreased accuracy and F1 score.
@@ 4.5 Effect of Beta on DPO and IPO
In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms. Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO). It controls the strength of the Kullback-Leibler (KL) divergence penalty between the trained model and the reference model within the loss function. Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
Similar to experiments in Section 4.3, we measured the average accuracy and F1 score across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in Figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75% and an average F1 score of 0.55 and 0.56. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3, and a loss of 0.13 in average F1 score. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (Figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increased, Mistral’s accuracy dropped from 57.7% to 36 %, and its F1 score decreased from 0.55 to 0.49, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %, and its F1 score decreased from 0.56 to 0.43. Our results corroborate the evidence from related studies [27], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
@@ 5 Conclusion
In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational objectives, aiming to foster optimal student learning. By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts. Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over the traditional supervised fine-tuning (SFT) approach, significantly improving the alignment of LLMs with pedagogical objectives. Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs. By curating a second preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions. This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.

### v3

@@ 4.3 Main Findings: SFT vs LHP
We observed substantial performance improvements across all three SFT models—Llama, Mistral, and Zephyr. Specifically, Llama achieved accuracy gains of 13.1%, 10.1%, and 10.9% from DPO, IPO, and KTO, respectively. Mistral demonstrated increases of 8.7%, 6.1%, and 7.4%, while Zephyr showed improvements of 50.0%, 50.2%, and 52.0% (Table 2). A similar trend was observed in F1 scores: Llama exhibited increases of 6.3%, 2.7%, and 3.2%, Mistral showed gains of 5.6%, 3.7%, and 4.3%, and Zephyr saw improvements of 29.3%, 26.7%, and 30.2% for DPO, IPO, and KTO, respectively. Notably, DPO and KTO significantly outperformed IPO across all models. Also note these results are for a fixed beta of $0.1$. As we show in section 4.6, beta hyperparameter tuning can further boost the performance of LHP algorithms.
This experiment shows the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models. All models – Llama, Mistral and Zephyr exhibit pronounced improvements using LHP algorithms compared to SFT.
4.4 Pedagogical Shifts: Perplexity Comparison of SFT and LHP
@@
Dataset Construction: The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5). By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1). With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4). This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart Action 2, and similarly for Actions 4 and 5.
We then use this newly constructed Action dataset to evaluate the perplexity of responses generated by the original instruct model (Base) to its SFT and LHP variants.
Figure 3: Comparison of multi-round performance for SFT vs. LHP methods (DPO, IPO, KTO) across Llama, Mistral, and Zephyr models. The graphs illustrate average accuracy over 8 conversation rounds, revealing the superior performance of LHP methods in maintaining pedagogical alignment across extended conversation context.
Figure 4: Performance of LHP algorithms (DPO, IPO, and KTO) as a function of beta. Our results indicate that KTO outperforms both DPO and IPO with optimal beta hyperparameter search.
Perplexity Findings: Our analysis, as shown in table 3), reveals interesting patterns across the three models: (A) Llama: The base model shows similar perplexity across all actions, indicating no strong preference for guidance or direct solutions. SFT increases perplexity for Actions 2 and 5, shifting towards more guidance. Both DPO and KTO further optimize this behavior, with DPO showing the lowest perplexity for Actions 1 and 4; (B) Mistral: The base model shows slightly higher perplexity for Actions 1 and 2. SFT dramatically increases perplexity for Actions 2 and 5, strongly favoring guidance. DPO and KTO further reduce perplexity for Actions 1 and 4, with DPO showing slightly better optimization; (C) Zephyr: The base model shows relatively uniform perplexity across actions. SFT significantly increases perplexity for Actions 2 and 5. DPO and KTO further optimize this behavior, with DPO showing lower perplexity for Actions 1 and 4, while KTO shows higher perplexity for Actions 2 and 5.
These results demonstrate that SFT partially shifts model behavior towards providing more guidance, but LHP methods (DPO and KTO) further optimize this behavior. For Mistral and Zephyr, LHP methods dramatically increase the perplexity gap between guidance actions (1 and 4) and direct solution actions (2 and 5), indicating a significant shift towards pedagogical alignment. The perplexity analysis provides compelling evidence for the effectiveness of LHP methods in achieving pedagogical alignment, consistently outperforming SFT across all three models.
@@
Llama (Base) 2.02 2.04 2.07 2.01
Llama (SFT) 1.77 2.44 1.87 2.45
Llama (DPO) 1.70 2.45 1.78 2.43
Llama (KTO) 1.74 2.48 1.81 2.46
Mistral (Base) 2.11 2.06 2.01 1.97
Mistral (SFT) 2.06 5.51 2.14 5.37
Mistral (DPO) 1.86 5.49 1.93 5.31
Mistral (KTO) 1.90 5.01 1.98 4.93
Zephyr (Base) 2.05 2.09 2.04 2.07
@@
Zephyr (DPO) 1.91 16.44 1.97 15.66
Zephyr (KTO) 1.97 16.76 2.03 16.16
Table 3: Perplexity analysis of model responses corresponding to different actions. Lower perplexity for Actions 1 and 4 indicates a higher likelihood of providing scaffolded guidance like hints, while higher perplexity for Actions 2 and 5 suggests a lower tendency to offer direct solutions. Results show SFT partially improves pedagogical alignment, while LHP methods (DPO, KTO) further optimize it by widening the perplexity gap between guidance (Actions 1, 4) and direct solution (Actions 2, 5) responses. For more details about actions please refer to Table 1.
4.5 Pedagogical Consistency Over Time: SFT vs. LHP in Extended Conversations
To assess the robustness and consistency of pedagogical alignment across extended conversations, we conducted an in-depth analysis of model accuracy across multiple conversation rounds. This analysis is important because as conversations progress, the context becomes more complex and lengthy, potentially challenging the model’s ability to maintain consistent pedagogical alignment.
The analysis reveals striking differences in pedagogical alignment between SFT and LHP methods across all three models. As can be seen in Figure 3, Llama’s SFT model shows a clear decline in alignment as conversations progress, dropping from $75\%$ to $60\%$ by round $8$, while LHP methods maintain consistent alignment around $75\%$. Mistral and Zephyr demonstrate even more pronounced contrasts. Their SFT models struggle with pedagogical alignment in extended conversations, with Mistral fluctuating between $63-69\%$ and Zephyr performing poorly at $11-30\%$. In stark contrast, all LHP approaches (DPO, IPO, KTO) for both models maintain high pedagogical alignment above $70\%$ throughout, with Zephyr showing the most dramatic improvement over its SFT counterpart. This consistent performance of LHP methods across all models, particularly in later conversation rounds, underscores their effectiveness in maintaining pedagogical alignment even as context complexity increases.
These results demonstrate that LHP-aligned models maintain high pedagogical alignment even with increasing context length, unlike SFT models which tend to degrade in performance over extended conversations. This robustness is crucial for effective tutoring, as it ensures the model can provide reliable guidance throughout a problem-solving session, regardless of its context length.
4.6 Effect of Beta on LHP Algorithms
@@
In this experiment, we analyze the impact of the beta parameter on the performance of LHP algorithms. Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO). It controls the strength of the Kullback-Leibler divergence penalty between the trained model and the reference model within the loss function. The choice of beta depends on the specific model and dataset and thus necessitates an empirical study. As we show next, beta hyperparameter tuning can significantly affect the performance of LHP algorithms, with KTO emerging as the top performer. Similar to experiments in Section 4.3, we measure the average accuracy and F1 score across the three classification fields. We find that the performance of all three algorithms is sensitive to the choice of beta, but to varying degrees and with different optimal values for each algorithm and model.
For DPO, we observe that lower beta values (0.1 or 0.3) generally yield the best results. Llama achieved the highest average accuracy of 77.3% and average F1 score of 0.57 with beta as 0.1. Mistral performed best with beta = 0.1, reaching 74.1% accuracy and 0.52 F1 score. Zephyr showed similar performance for beta = 0.1 and 0.3, with a slight edge for 0.1 (73.4% accuracy, 0.55 F1 score).
IPO demonstrated less sensitivity to beta changes, especially for Llama and Mistral. Llama’s performance peaked at beta = 0.6 with 74.9% accuracy and 0.54 F1 score. Mistral showed minimal variation, with slightly better results at beta = 0.6 (71.4% accuracy, 0.51 F1 score). Zephyr performed best with beta = 0.3, achieving 72.6% accuracy and 0.55 F1 score.
KTO emerged as the top performer, showing consistent improvement with increasing beta values across all models. For Llama, performance peaked at beta = 0.9 with 78.1% accuracy and 0.57 F1 score. Mistral also showed best results at beta = 0.9, reaching 74.7% accuracy and 0.51 F1 score. Zephyr demonstrated steady improvement, with optimal performance at beta = 0.9 (75.7% accuracy, 0.55 F1 score).
Notably, with optimal beta values, KTO outperforms both DPO and IPO across all models, emphasizing its effectiveness in pedagogical alignment tasks when properly tuned. Our findings underscore the necessity of careful hyperparameter tuning in LHP algorithms to achieve optimal pedagogical alignment. They also suggest that KTO, when properly tuned, may be particularly well-suited for tasks requiring sustained pedagogical alignment.
@@ 5 Conclusion
In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational goals, aiming to foster optimal student learning outcomes. By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts. Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over standard supervised fine-tuning (SFT), significantly improving the alignment of LLMs with pedagogical objectives. Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs. By curating a preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions. This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.
6 Limitations
While our methodology shows promising results, there are certain inherent limitations that are challenging to overcome. One such limitation is the dynamic and complex nature of the educational landscape. The effectiveness of our approach may vary depending on a multitude of factors, including the diversity of student learning styles, the evolving nature of educational content, and the rapid advancements in AI technology. Additionally, computational constraints may also limit the scalability of our approach in larger, more diverse educational settings. More comprehensive user studies can assess the impact of these aligned models in real-world educational settings.
7 Ethics and Risks
The use of LLMs in education comes with ethical considerations and potential risks. These include the need to ensure the models do not amplify biases, the importance of protecting sensitive student data, and the necessity of maintaining human guidance in the learning process. In addition, the impact of LLMs on learning outcomes should be continuously evaluated to ensure their effectiveness and avoid unintended negative consequences.
Acknowledgments
This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.